O GEOMETRY Surface area of a cylinder the surface area of a cylinder with a height of 6 m and a base radius of 2 m. the valuehelp solving please
Answer #1

Height h = 6 m

Radius r = 2 m

Surface Area of the cylinder

= 2\pi rh + 2\pi r^2

= 2(3.14)(2)(6) + 2(3.14) (2)^2

= 75.36 + 25.12

= 100.48 m^2

Therefore,

Answer: 100.48 m^2